The sum of three consecutive integers is 369. What is the smallest integer?
Answer by askmemath(368)

If they are consecutive we can write the numbers as n,n+1,n+2
So n+1+n+2+n = 369
3n+3 = 369
3n = 366 Subtracting 3 on both sides
n = 122 Dividing by 3 throughout
n+1 = 123
n+2 = 124

Check:
122+123+124 = 369 Yay!!